OpenStudy (anonymous):

the speed of the A&M freight train is 14 mph less than the speed of the A&M passanger train.the passanger train travels 400mi in the same time that thw freight train travels 330mi Find the speed of each train. thanks

OpenStudy (anonymous):

system of linear equation application. . .

hero (hero):

d = rt freight train: 330 = (r-14)t passenger train 400 = rt 330 = rt - 14t 400 = rt 330 = 400 - 14t 14t = 400 - 330 14t = 70 t = 5 r = 80

hero (hero):

r - 14 = 80 - 14 = 66 Freight = 66 Passenger = 80

OpenStudy (anonymous):

wow, thanks Hero ;)

hero (hero):

That's the smoothest solution you'll ever see